OpenStudy (anonymous):

[3.06] Given the function f(x) = −2x + 8, find x if f(x) = 22. −36 −7 52 −15

undeadknight26 (undeadknight26):

easy: f(x) = 2x + 8 f(22) = 2 x 22 +8 f(22) = 44 + 8 f(22) = 52

OpenStudy (anonymous):

oooooooo I know the steps now

OpenStudy (anonymous):

one more

undeadknight26 (undeadknight26):

aye :D

undeadknight26 (undeadknight26):

kk but u try first.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let A = {7, 8, 9, 10} and let B = {3, 4}. Find A X B. {3, 4, 7, 8, 9, 10} { } {(7, 3), (7, 4), (8, 3), (8, 4), (9, 3), (9, 4), (10, 3), (10, 4)} {3, 4}

OpenStudy (anonymous):

what is this

OpenStudy (anonymous):

its hard

undeadknight26 (undeadknight26):

i have no idea...

OpenStudy (anonymous):

hold on I think its b

OpenStudy (anonymous):

if your finding x you would want to make 22=f(x) it would be 22=-2x+8 if it gave you x and you wanted to find f(x) that is when you would plug it in for x

OpenStudy (anonymous):

yeah so its b

OpenStudy (anonymous):

I don't know about that set problem. I haven't worked with a problem like that since highschool